OpenStudy (anonymous):

What is the mass of 3.5 moles of NaCl? A. 5.8 × 101 grams B. 3.5 × 100 grams C. 2.0 × 102 grams D. 6.2 × 101 grams E. 6.0 × 10–2 grams

thomaster (thomaster):

Look at this. what do you have to do to solve this?

thomaster (thomaster):

You still there @RH ?

thomaster (thomaster):

Look on the image i attached, where you go from mol to mass, it says multiply with molar mass So multiply 3.5 with the molar mass of NaCl and you'll have your answer

OpenStudy (mayankdevnani):

What is the formula mass of NaCl. @RH

OpenStudy (anonymous):

Yes I am here, sorry my laptop frooze

OpenStudy (mayankdevnani):

so can you find out the formula mass of NaCl ?

OpenStudy (mayankdevnani):

@RH can you?

OpenStudy (anonymous):

yes, it is 58.4425

OpenStudy (anonymous):

is C the correct answer?

OpenStudy (mayankdevnani):

lets assume that formula mass of NaCl be 58.5 u .

OpenStudy (mayankdevnani):

okk ???

OpenStudy (anonymous):

ok

OpenStudy (mayankdevnani):

so, \[\huge \bf mole=\frac{weight/mass}{formula~mass}\] so, \[\huge \bf 3.5=\frac{Mass}{58.5}\] now cross multiply them, \[\huge \bf Mass=3.5 \times 58.5\]

OpenStudy (mayankdevnani):

Can you find out the value of mass? @RH

OpenStudy (chmvijay):

1 mole of Nacl =58.5 gram of Nacl 3.5 moles of Nacl = 58.5* 3.5 = ?

OpenStudy (mayankdevnani):

@RH are you still here?

OpenStudy (anonymous):

204.75

OpenStudy (mayankdevnani):

correct

OpenStudy (anonymous):

so C is the correct answer?

OpenStudy (mayankdevnani):

yeah.

OpenStudy (anonymous):

thanks:)

OpenStudy (mayankdevnani):

welcome :)

OpenStudy (chmvijay):

:)